How many grams of NaCl is required prepare 200 mL of a solution of 1 M NaCl (Molar of mass of NaCl=58.44g/mol)?


1
Expert's answer
2021-09-27T01:48:40-0400

200 ml of 1 M Nacl have

Moles of NaCl = 200 × 1 / 1000 = 0.2 mole

Mass of NaCl = moles of NaCl × Molar mass .

Mass of NaCl = 0.2 × 58.44 = 11.688 gram .